OpenStudy (anonymous):

verify the following identity sin2t-cot t=-cot t*cos2t

OpenStudy (campbell_st):

just a quick question... is it 1. sin(2t) or 2. \[\sin^2(t)\]

OpenStudy (anonymous):

@campbell_st Sin(2t)

OpenStudy (anonymous):

option 1

OpenStudy (campbell_st):

ok... start with the left hand side \[\sin(2t) = 2\sin(t) \cos(t)\] and \[\cot(t) = \frac{\cos(t)}{\sin(t)}\] so you have \[2\sin(t)\cos(t) - \frac{\cos(t)}{\sin(t)}\] you noe need a common demoninator so multiply the 1st term by sin(t) and you get \[\frac{2\sin^2(t) \cos(t) - \cos(t)}{\sin(t)} = \frac{\cos(t)(2\sin(t) - 1)}{\sin(t)}\] which is \[\cot(t) \times (2\sin^2(t) - 1)\] now look at the different forms of \[\cos(2t) \] and I think you'll find there is one where \[\cos(2t) = 1 - 2\sin^2(t) \] so reversing the order means you'll get -cos(2t) and hopefully its obvious from here